OpenStudy (abbles):

Prove the following identities:

OpenStudy (abbles):

sin(x-y)sin(x+y) = sin^2x - sin^2y

OpenStudy (abbles):

Here's what I have so far (not sure if it's right...) sin(x-y)sin(x+y) = (sinxcosy - cosxsiny)(sinxcosy + cosxsiny) (sinxcosy)^2 - (cosxsiny)^2 Then I'm thinking of using the product to sum identities... how am I doing so far? It seems a bit too complicated, but I dunno.

OpenStudy (crisulcampo):

[sin(x)cos(y) - cos(x)sin(y)][sin(x)cos(y)] + cos(x)sin(y)] = sin^2 (x) - sin^2 (y) you should know that (a-b)(a+b) = a^2 - b^2 so: [sin(x)cos(y)]^2 - [cos(x)sin(y)]^2 = sin^2 (x) - sin^2 (y) sin^2 (x) [ 1 - sin^2 (y)] - sin^2 (y) [1- sin^2 (x)] = sin^2 (x) - sin^2 (y) sin^2 (x) - sin^2 (y)sin^2(x) - sin^2(y) + sin^2 (y)sin^2(x) = sin^2 (x) - sin^2 (y) final result is: sin^2(x) - sin^2 (y) = sin^2 (x) - sin^2 (y) I hope you understand

OpenStudy (abbles):

Thank you so much! That is very helpful, I'll go through it. Do you mind looking over the steps I tried and seeing if I was doing it correctly or not? I posted what I had so far above.

zepdrix (zepdrix):

From this point,\[\large\rm (\sin x \cos y)^2 - (\cos x \sin y)^2\]\[\large\rm \sin^2x \color{orangered}{\cos^2y}-\color{orangered}{\cos^2x}\sin^2y\]you need to get rid of all the cosines, ya? So I would apply your `Pythagorean Identity` to these orange parts.

OpenStudy (crisulcampo):

what you did it's OK.... if you have a doubt about my solution method, let me know please my friend

OpenStudy (abbles):

sin^2x(1 - sin^2y) - (1 - cos^2x)(sin^2y) Like this zep?

OpenStudy (abbles):

Whoops, that should be (1 - sin^2x) in the second part ;)

zepdrix (zepdrix):

ya ya ya, good stuff

zepdrix (zepdrix):

and mooooore distributing D: stuff should cancel out.

OpenStudy (abbles):

sin^2x(1 - sin^2y) - (1 - sin^2x)(sin^2y) sin^2x - sin^2y(sin^2x) - (sin^2y - sin^2x(sin^2y)) It looks like a mess :P But do you know what to do from here?

OpenStudy (crisulcampo):

yes

OpenStudy (crisulcampo):

abbles, I am providing you of a word archive with the specific points, every single step well explained do you agree?

OpenStudy (abbles):

hmmm.. what?

OpenStudy (crisulcampo):

wait a second, I'm starting solving from your post

OpenStudy (crisulcampo):

what you need to do from there is just apply distributive property, then you'll note there is a possitive\[\sin^2 (x)\sin^2 (y)\] and a negative one, so just cancel them out

OpenStudy (abbles):

So like this? sin^2x - sin^2ysin^2x - sin^2y + sin^2xsin^2y I see what you mean! I would be left with sin^2x - sin^2y Exactly like the original :) Thanks!

zepdrix (zepdrix):

yay!

OpenStudy (crisulcampo):

dare yourself! try to get the left expression starting from sin^2 (y) - sin^2(x) at the beginning... are you able to do it, abbles?